# The Analysis of Willmore Surfaces and its Generalizations in Higher Dimensions

### v1

@@ I Introduction
The study of topological and geometrical properties of manifolds through the search of special metrics “equipping” these manifolds and solving some special Partial Differential Equation has a long and rich history which takes its origin maybe in the construction of special curves (like Brachistochrone curves in the XVIIth century, or Euler’s elastica in the XVIIIth century) and constant Gauss curvature metrics on closed surfaces in relation with the uniformization theorem for Riemann Surfaces in the XVIIIth century. The development of what could be called “intrinsic geometric analysis” has been the source of spectacular results in differential topology, differential geometry and complex geometry with the search of constant scalar curvature metric, Einstein metrics, Kähler Einstein metrics, solutions to the Ricci flow… Another branch of geometric analysis is dealing with the study of “special submanifolds” within a given Riemannian manifold and its interaction with the geometry of the manifold itself. This branch of geometric analysis is maybe rooted originally both in the calculus of variations with the variational constructions of closed geodesics, the resolution of the Lagrange-Plateau problem in Euclidean space as well in the “explicit differential geometry” of special submanifolds such as plane algebraic curves, algebraic surfaces, etc. The central objects of what can be described as “extrinsic geometric analysis” are minimal surfaces and their generalizations (constant and prescribed mean curvature surfaces). The underlying PDEs are related to the area variations under various constraints. In the present work, we consider a branch of extrinsic variational geometric analysis that has seen substantial development in recent decades, and concerns immersions that arise as critical points of Lagrangians depending on the first and second fundamental form, and their covariant derivatives. The simplest and maybe most studied model for such a Lagrangian is the Willmore functional of immersed surfaces into ${\mathbb{R}}^{m}$. Let $\Sigma$ be closed oriented two dimensional manifold and let $\vec{\Phi}$ be a $C^{2}$ immersion of this surface into the three dimensional Euclidean space ${\mathbb{R}}^{m}$. Denoting by $g_{\vec{\Phi}}$ the first fundamental form of this immersion and by $\vec{\mathrm{I\!I}}_{\vec{\Phi}}$ its second fundamental form, the mean curvature of the immersion is given by
@@
This energy has been initially introduced by Sophie Germain and Simeon Poisson in an attempt to generalize to two dimensional elastic membranes the famous Euler Elastica modelizing the free energy of a beam [64, 160]. The very first derivation of the Euler–Lagrange equation of $W$ was made by Poisson himself around 1814 in the case of a graph (see page 60 of [160] and chapter 10 of [137]) in $\mathbb{R}^{3}$. Using concepts which were not completely clarified at the time (such as the Gauss curvature and the Laplace–Beltrami operator), we can rewrite the Poisson’s Euler–Lagrange equation for the Lagrangian $W$ in the following form
@@
where $H_{\vec{\Phi}}=\vec{n}_{\vec{\Phi}}\cdot\vec{H}_{\vec{\Phi}}$ is the mean curvature and $\vec{n}_{\vec{\Phi}}$ the unit Gauss map to the immersion, $\Delta_{g}$ is the negative Laplace–Beltrami operator which reads in local coordinates $(x^{1},x^{2})$ $\Delta_{g}\cdot\coloneqq(\det(g))^{-1/2}\partial_{x_{i}}(g^{ij}\,(\det(g))^{1/2}\partial_{x_{j}}\cdot)$ and $K_{\vec{\Phi}}\coloneqq\mbox{det}_{g_{\vec{\Phi}}}(\vec{n}_{\vec{\Phi}}\cdot\vec{\mathrm{I\!I}}_{\vec{\Phi}})$ is the Gauss curvature. Throughout this paper, we shall often write $g=g_{\vec{\Phi}}$ when there is no ambiguity. With these notations we have also
@@
Comparing (I.2) with (I.1), it appears that Poisson’s Euler–Lagrange equation is a degenerate nonlinear fourth-order equation combining analytical difficulties which were way too advanced for the early XIXth century. Hence in a natural way, in the following 100 years, the attention has been exclusively devoted to special solutions solving a simpler second order version of (I.1):
@@
Poisson equation (I.1) has been rediscovered by Wilhelm Balschke [21] and his student Gerhard Thomsen [183] a century after (around 1924) in the framework of conformal geometry. Hence, the study of the Lagrangian $W$ was motivated by the merging of two theories: conformal geometry and minimal surfaces theory. On the one hand they proved that $W$ is invariant under the action of generic [^1] Precisely, for any conformal transformation $\Psi$ from ${\mathbb{R}}^{3}\cup\{\infty\}$ into itself and for any immersion $\vec{\Phi}$ of an oriented closed two dimensional manifold $\Sigma$ into ${\mathbb{R}}^{3}$ such that $\Psi^{-1}(\{\infty\})\cap\vec{\Phi}(\Sigma)=\emptyset$ there holds $W(\Psi\circ\vec{\Phi})=W(\vec{\Phi})\ .$ conformal transformations on the other hand, as it has been observed above, minimal surfaces are stable critical points of $W$. Probably because of the absence of known examples that were not just composition of minimal immersions with conformal transformations, the interest in functional $W$ disappeared again from the mathematical literature for a few decades until the famous paper [188] by Thomas Willmore in 1965. In this paper Willmore proved that for any immersion $\vec{\Phi}$ of a closed surface $\Sigma$ the following lower bound holds
@@
The first analytical work on the Willmore functional $W$ is the paper by Leon Simon [180] in which he shows that the infimum of $W$ is achieved among all possible $C^{2}$ immersions of the torus by a smooth immersion. Relying for a large part on the analysis of [180], Matthias Bauer and Ernst Kuwert in [12], by proving some clever strict inequalities on connected sums of immersions of surfaces, have been able to extend Simon’s result to any genus, that is, for any genus $g$ the infimum of $W$ among all immersions of the surface of genus $g$ is achieved. What should be noticed regarding the analysis in [180] is that first the minimization is considered in measure theoretical sense using the theory of varifolds, then the regularity is relying a lot on comparison arguments with local competitors which are graphs satisfying the biharmonic equation (i.e. the linearised version of (I.1)). In this sense, the analysis in [180] is very much restricted to minimization operations.
The third author of the present work has been considering in the early 2000 the possibility to use more classical functional analysis to perform the minimisation of $W$ with the perspective of considering more general variational arguments such as min-max operations, or the study of the associated gradient flow. One of the first difficulties was to obtain an Euler–Lagrange equation compatible with the functional space in which the variations of $W$ are studied. In order to illustrate this difficulty, it could be interesting to look at a simpler framework and to go one dimension lower. We consider the variations of the Euler Elastica for immersions $\vec{\Phi}$ of the segment $[0,1]$ into the plane $\mathbb{R}^{2}$ given by
@@
But this is not the solvability of (I.3) which is addressed at this stage, but rather the compatibility between the Euler–Lagrange equation and the Lagrangian from a purely variational perspective. In unit speed parameterization $\vec{\Phi}(t)$ for $s\in[0,L]$ and $L=\int dl_{\vec{\Phi}}$ we have $\kappa_{\vec{\Phi}}\,\vec{n}_{\vec{\Phi}}=\frac{d^{2}\vec{\Phi}}{ds^{2}}$ and thus it holds
@@
This does not define a distribution for an arbitrary $\vec{\Phi}$ in $W^{2,2}([0,L],{\mathbb{R}}^{2})$. As paradoxal as it appears, the Euler–Lagrange equation (I.3) is not compatible with the Lagrangian itself!
@@
$2\,H_{\vec{\Phi}}\,|H_{\vec{\Phi}}|^{3}\qquad\notin L^{1}_{\textup{loc}}(\Sigma).$
@@
The Gauss–Bonnet theorem holds, denoting $\gamma(\Sigma)$ is the genus of $\Sigma$,
@@
A proof of the first part of this result was proposed in [168] but a very last ingredient which requires $\vec{\Phi}\in W^{2,2}(\Sigma,{\mathbb{R}}^{m})$ and not only $\vec{n}_{\vec{\Phi}}\in W^{1,2}(\Sigma,G_{2}({\mathbb{R}}^{m}))$ was missing to complete the argument and to ensure that $\phi$ given by (1.72) in the middle of page 323 is a bi-Lipschitz homeomorphism. The complete argument for proving Theorem I.1 is given in Section IV of the present paper. Observe that the existence of bi-Lipschitz isothermal charts does not hold if one only assumes that the Gauss map is in $W^{1,2}$ without assuming that the weak immersion is in $W^{2,2}$: in the local sense, a counterexample $z\rightarrow(z^{2}/|z|,0)\in\mathbb{R}^{3}$ is provided by the first author in [104], and when $\Sigma=\mathbb{S}^{2}$, Plotnikov [159] construct a global counterexample by composing the map $z\rightarrow z^{2}/|z|$ with the stereographic projection.
@@
Some weeks after the proof of Theorem I.2 has been posted on arxiv, a proof of this result has appeared on arxiv as well but assuming the sequence $\vec{\Phi}_{k}$ to be conformal with respect to some sequence of Riemann surface (see [100]). It is important to insist at this stage that the restriction to conformal weak immersions is missing the goal posed in [166] of developing a variational theory for Willmore. It does not fulfill the need of exploring the neighborhood of a weak immersion in order to deduce an Euler–Lagrange equation for instance and Theorem I.1 is an essential tool in that respect.
The notion of branched $W^{2,2}$ weak immersion into ${\mathbb{R}}^{m}$ is an extension of $W^{2,2}$ weak immersion allowing for isolated branched points $\{a_{1},\dots,a_{L}\}$ at which the gradient of the weak immersion is degenerating. Branched $W^{2,2}$ weak immersion have an underlying conformal structure that is there exists a constant Gauss curvature metric $h$ on $\Sigma$ such that
@@
But the conformal factor $\alpha$ is not bounded in $L^{\infty}$ anymore at each branched point but has a relatively well understood asymptotic expansion: in a local conformal chart $x=(x^{1},x^{2})\in D^{2}$ for $h$ containing exactly one of the branched point, say $a_{i}$, and for which $x(a_{i})=0$, there exists $d_{i}\in{\mathbb{N}}$ such that (see Lemma A.5 of [166])
$\|\alpha(x)-d_{i}\log|x|\|_{L^{\infty}(D^{2})}<+\infty,$
@@
$2\vec{H}=\nabla S\cdot\nabla^{\perp}\vec{\Phi}+\nabla\vec{R}\,\dot{\times}\,\nabla^{\perp}\vec{\Phi}.$ (I.14)
While the use of local conformal coordinates was centrally used in [164], Section V of the present paper is presenting a proof of the regularity without using conformal coordinates. The motivation being to “prepare the ground” for considering regularity questions for solution to the Euler–Lagrange equation to higher dimensional counterpart to the Willmore Lagrangian which is the subject of the second part of the paper. Before coming to this second part we prove the following approximation theorem (in Section IV.4.3), which is one of the main new result of the present paper.
@@
In 2018, Andrea Mondino and Huy Nguyen [146] proved that the Willmore functional is the only (up to a topological constant) conformally invariant functional among all the functionals defined on hypersurfaces of 3-dimensional manifolds and depending only on the first and second fundamental form. However, this is not the case in higher dimensions and codimensions. Easy examples can be found in dimension 4, with for instance, functionals defined on immersed submanifolds $\Sigma^{4}\subset\mathbb{R}^{5}$ such as
@@
The volume renormalization procedure for minimal hypersurfaces has the advantage of producing conformally invariant functionals having minimal hypersurfaces as critical points. However, as discussed in Section II.2.1, some of these functionals are not bounded from below or from above. Returning to the original idea of Germain and Poisson [64, 160], one would like to study functionals that measure in some sense the umbilicity of a hypersurface. Such a functional should be able to provide ”best representation” of submanifolds and in particular, this requires to have a lower bound. In [70], Gover–Waldron proved that conformally invariant functionals can also be stated in terms of singular Yamabe-type problem, a problem raised by Loewner–Nirenberg [113] in 1974. In [67, 70], the authors prove that given any separating submanifold $\Sigma^{d-1}$ of some Riemannian manifold $(M^{d},g)$, if one look for the ”best” asymptotic expansion of a defining function [^3] A defining function $s\in C^{\infty}(M)$ is defining for $\Sigma$ if $s=0$ and $|ds|_{g}=1$ on $\Sigma$. The question raised in [70] is to find the highest number $\ell\in\mathbb{N}$ and a defining function $s$ such that we have an expansion of the form $|ds|_{g}=1+O(s^{\ell})$ near $\Sigma$. $s$ for $\Sigma$, then there is an obstruction, called the obstruction density. This obstruction density for d1=2d-1=2 (i.e. when $\Sigma$ is a surface) and $M^{d}=\mathbb{R}^{3}$ turns out to be exactly the left-hand side of (I.1). As proved by [72, 68], this obstruction can be understood in general as the Euler–Lagrange equation of a coefficient in some renormalized volume expansion reminiscent to [75, 74]. These alternatives will be discussed in Section II.2.2. One common point of these functionals is that if $\Sigma^{n}\subset\mathbb{R}^{d}$ has even dimension $n\geq 4$, then the Euler–Lagrange equation of any of this generalized Willmore energy has a leading-order term of the form $\Delta_{g_{\Sigma}}^{\frac{n}{2}}\vec{H}_{\Sigma}$. This term is also the leading-order term of the functional
$\vec{\Phi}\in\mathrm{Imm}(\Sigma^{n};\mathbb{R}^{d})\mapsto\int_{\Sigma^{n}}\big|\nabla^{\frac{n-2}{2}}\vec{\mathrm{I\!I}}_{\vec{\Phi}}\big|^{2}_{g_{\vec{\Phi}}}\ d\textup{vol}_{g_{\vec{\Phi}}}.$

### v3

@@ I Introduction
The study of topological and geometrical properties of manifolds through the search of special metrics “equipping” these manifolds and solving some special Partial Differential Equations has a long and rich history which takes its origin maybe in the construction of special curves (like Brachistochrone curves in the XVIIth century, or Euler’s elastica in the XVIIIth century) and constant Gauss curvature metrics on closed surfaces in relation with the uniformization theorem for Riemann Surfaces in the XXth century. The development of what could be called “intrinsic geometric analysis” has been the source of spectacular results in differential topology, differential geometry, and complex geometry with the search of constant scalar curvature metrics, Einstein metrics, Kähler–Einstein metrics, solutions to the Ricci flow… Another branch of geometric analysis is dealing with the study of “special submanifolds” within a given Riemannian manifold and its interaction with the geometry of the manifold itself. This branch of geometric analysis is maybe rooted originally both in the calculus of variations with the variational constructions of closed geodesics, the resolution of the Lagrange-Plateau problem in Euclidean space as well in the “explicit differential geometry” of special submanifolds such as plane algebraic curves, algebraic surfaces, etc. The central objects of what can be described as “extrinsic geometric analysis” are minimal surfaces and their generalizations (constant and prescribed mean curvature surfaces). The underlying PDEs are related to the area variations under various constraints. In the present work, we consider a branch of extrinsic variational geometric analysis that has seen substantial development in recent decades, and concerns immersions that arise as critical points of Lagrangians depending on the first and second fundamental form, and their covariant derivatives. The simplest and maybe most studied model for such a Lagrangian is the Willmore functional of immersed surfaces into ${\mathbb{R}}^{m}$ with $m\geq 3$. Let $\Sigma$ be a closed oriented two-dimensional manifold and let $\vec{\Phi}$ be a $C^{2}$ immersion of this surface into ${\mathbb{R}}^{m}$. Denoting by $g_{\vec{\Phi}}$ the first fundamental form of this immersion and by $\vec{\mathrm{I\!I}}_{\vec{\Phi}}$ its second fundamental form, the mean curvature of the immersion is given by
@@
This energy has been initially introduced by Sophie Germain and Siméon Denis Poisson in an attempt to generalize to two-dimensional elastic membranes the famous Euler Elastica modeling the free energy of a beam [64, 163]. The very first derivation of the Euler–Lagrange equation of $W$ was made by Poisson himself around 1814 in the case of a graph (see page 60 of [163] and chapter 10 of [141]) in $\mathbb{R}^{3}$. Using concepts which were not completely clarified at the time (such as the Gauss curvature and the Laplace–Beltrami operator), we can rewrite Poisson’s Euler–Lagrange equation for the Lagrangian $W$ in the following form
@@
where $H_{\vec{\Phi}}=\vec{n}_{\vec{\Phi}}\cdot\vec{H}_{\vec{\Phi}}$ is the mean curvature and $\vec{n}_{\vec{\Phi}}$ the unit Gauss map to the immersion, $\Delta_{g}$ is the negative Laplace–Beltrami operator which reads in local coordinates $(x^{1},x^{2})$ $\Delta_{g}\cdot\coloneqq(\det(g))^{-1/2}\partial_{x_{i}}(g^{ij}\,(\det(g))^{1/2}\partial_{x_{j}}\cdot)$ and $K_{\vec{\Phi}}\coloneqq\mbox{det}_{g_{\vec{\Phi}}}(\vec{n}_{\vec{\Phi}}\cdot\vec{\mathrm{I\!I}}_{\vec{\Phi}})$ is the Gauss curvature. [^1] The Euler-Lagrange equation in higher codimension has the same cubic structure in the second fundamental form, see [190, Theorem 2.1]. Throughout this paper, we shall often write $g=g_{\vec{\Phi}}$ when there is no ambiguity. With these notations we have also
@@
Comparing (I.2) with (I.1), it appears that Poisson’s Euler–Lagrange equation is a degenerate nonlinear fourth-order equation combining analytical difficulties which were way too advanced for the early XIXth century. Hence in a natural way, in the following century, the attention has been exclusively devoted to special solutions solving a simpler second order version of (I.1):
@@
Poisson’s equation (I.1) was rediscovered by W. Schadow [^2] See the comment by Wilhelm Blaschke in [21, Exercise 7, $\S$ 83]. and also appeared in the PhD work of Gerhard Thomsen [186] a century later (in 1923) in the framework of conformal geometry. Hence, the study of the Lagrangian $W$ was motivated by the merging of two theories: conformal geometry and minimal surfaces theory. On the one hand they proved that $W$ is invariant under the action of generic [^3] Precisely, for any conformal transformation $\Psi$ from ${\mathbb{R}}^{3}\cup\{\infty\}$ into itself and for any immersion $\vec{\Phi}$ of an oriented closed two dimensional manifold $\Sigma$ into ${\mathbb{R}}^{3}$ such that $\Psi^{-1}(\{\infty\})\cap\vec{\Phi}(\Sigma)=\emptyset$ there holds $W(\Psi\circ\vec{\Phi})=W(\vec{\Phi})\ .$ conformal transformations on the other hand, as it has been observed above, minimal surfaces are stable critical points of $W$. Probably because of the absence of known examples that were not just composition of minimal immersions with conformal transformations, the interest in functional $W$ disappeared again from the mathematical literature for a few decades until the famous paper [192] by Thomas Willmore in 1965. In this paper Willmore proved that for any immersion $\vec{\Phi}$ of a closed surface $\Sigma$ the following lower bound holds
@@
The first analytical work on the Willmore functional $W$ is the paper by Leon Simon [183] in which he shows that the infimum of $W$ is achieved among all possible $C^{2}$ immersions of the torus by a smooth immersion. Relying for a large part on the analysis of [183], Matthias Bauer and Ernst Kuwert in [12], by proving some clever strict inequalities on connected sums of immersions of surfaces, have been able to extend Simon’s result to any genus, that is, for any genus $g$ the infimum of $W$ among all immersions of the surface of genus $g$ is achieved. What should be noticed regarding the analysis in [183] is that first the minimization is considered in a measure-theoretical sense using the theory of varifolds, then the regularity relies a lot on comparison arguments with local competitors which are graphs satisfying the biharmonic equation (i.e. the linearised version of (I.1)). In this sense, the analysis in [183] is very much restricted to minimization operations.
The third author of the present work has been considering in the early 2000s the possibility to use more classical functional analysis to perform the minimization of $W$ with the perspective of considering more general variational arguments such as min-max operations, or the study of the associated gradient flow. One of the first difficulties was to obtain an Euler–Lagrange equation compatible with the functional space in which the variations of $W$ are studied. In order to illustrate this difficulty, it could be interesting to look at a simpler framework and to go one dimension lower. We consider the variations of the Euler Elastica for immersions $\vec{\Phi}$ of the segment $[0,1]$ into the plane $\mathbb{R}^{2}$ given by
@@
However this is not the solvability of (I.3) which is addressed at this stage, but rather the compatibility between the Euler–Lagrange equation and the Lagrangian from a purely variational perspective. In unit speed parameterization $\vec{\Phi}(t)$ for $s\in[0,L]$ and $L=\int dl_{\vec{\Phi}}$ we have $\kappa_{\vec{\Phi}}\,\vec{n}_{\vec{\Phi}}=\frac{d^{2}\vec{\Phi}}{ds^{2}}$ and thus it holds
@@
This does not define a distribution for an arbitrary $\vec{\Phi}$ in $W^{2,2}([0,L],{\mathbb{R}}^{2})$. As paradoxical as it appears, the Euler–Lagrange equation (I.3) is not compatible with the Lagrangian itself!
@@
$2\,H_{\vec{\Phi}}\,|H_{\vec{\Phi}}|^{2}\qquad\notin L^{1}_{\textup{loc}}(\Sigma).$
@@
The Gauss–Bonnet theorem holds, denoting by $\gamma(\Sigma)$ the genus of $\Sigma$,
@@
A proof of the first part of this result was proposed in [171] but a very last ingredient which requires $\vec{\Phi}\in W^{2,2}(\Sigma,{\mathbb{R}}^{m})$ and not only $\vec{n}_{\vec{\Phi}}\in W^{1,2}(\Sigma,G_{2}({\mathbb{R}}^{m}))$ was missing to complete the argument and to ensure that $\phi$ given by [171, Equation (1.72)] in the middle of page 323 was a bi-Lipschitz homeomorphism. The complete argument for proving Theorem I.1 is given in Section IV of the present paper. Observe that the existence of bi-Lipschitz isothermal charts does not hold if one only assumes that the Gauss map is in $W^{1,2}$ without assuming that the weak immersion is in $W^{2,2}$: in the local sense, a counterexample $z\rightarrow(z^{2}/|z|,0)\in\mathbb{R}^{3}$ is provided by the first author in [107], and when $\Sigma=\mathbb{S}^{2}$, Plotnikov [162] construct a global counterexample by composing the map $z\rightarrow z^{2}/|z|$ with the stereographic projection.
@@
Some weeks after the proof of Theorem I.2 was posted on arXiv, a proof of this result appeared on arXiv as well but assuming the sequence $\vec{\Phi}_{k}$ to be conformal with respect to some sequence of Riemann surface (see [103]). It is important to insist at this stage that the restriction to conformal weak immersions is missing the goal posed in [169] of developing a variational theory for the Willmore energy. It does not fulfill the need of exploring the neighborhood of a weak immersion in order to deduce an Euler–Lagrange equation for instance and Theorem I.1 is an essential tool in that respect.
The notion of branched $W^{2,2}$ weak immersion into ${\mathbb{R}}^{m}$ is an extension of $W^{2,2}$ weak immersion allowing for isolated branched points $\{a_{1},\dots,a_{L}\}$ at which the gradient of the weak immersion is degenerating. Branched $W^{2,2}$ weak immersions have an underlying conformal structure, that is to say, there exists a constant Gauss curvature metric $h$ on $\Sigma$ such that
@@
However the conformal factor $\alpha$ is not bounded in $L^{\infty}$ anymore at each branched point but has a relatively well understood asymptotic expansion: in a local conformal chart $x=(x^{1},x^{2})\in D^{2}$ for $h$ containing exactly one of the branched point, say $a_{i}$, and for which $x(a_{i})=0$, there exists $d_{i}\in{\mathbb{N}}$ such that (see Lemma A.5 of [169])
$\|\alpha(x)-d_{i}\log|x|\|_{L^{\infty}(D^{2})}<+\infty.$
@@
$2\vec{H}=\nabla S\cdot\nabla^{\perp}\vec{\Phi}+\nabla\vec{R}\times\nabla^{\perp}\vec{\Phi}.$ (I.14)
While the use of local conformal coordinates was essential in [167], Section V of the present paper is presenting a proof of the regularity without using conformal coordinates. The motivation is to “prepare the ground” for considering regularity questions for solution to the Euler–Lagrange equation to higher dimensional counterpart to the Willmore Lagrangian, which is the subject of the second part of the paper. Before coming to this second part we prove the following approximation theorem (in Section IV.4.3), which is one of the main new result of the present paper.
@@
In 2018, Andrea Mondino and Huy The Nguyen [149] proved that the Willmore functional is the only (up to a topological constant) conformally invariant functional among all the functionals defined on hypersurfaces of 3-dimensional manifolds and depending only on the first and second fundamental form. However, this is not the case in higher dimensions and codimensions. Easy examples can be found in dimension 4, with for instance, functionals defined on immersed submanifolds $\Sigma^{4}\subset\mathbb{R}^{5}$ such as
@@
The volume renormalization procedure for minimal hypersurfaces has the advantage of producing conformally invariant functionals having minimal hypersurfaces as critical points. However, as discussed in Section II.2.1, some of these functionals are not bounded from below or from above. Returning to the original idea of Germain and Poisson [64, 163], one would like to study functionals that measure in some sense the umbilicity of a hypersurface. Such a functional should be able to provide “best representation” of submanifolds and in particular, this requires to have a lower bound. In [70], Gover–Waldron proved that conformally invariant functionals can also be stated in terms of singular Yamabe-type problem, a problem raised by Loewner–Nirenberg [116] in 1974. In [67, 70], the authors prove that given any separating submanifold $\Sigma^{d-1}$ of some Riemannian manifold $(M^{d},g)$, if one looks for the “best” asymptotic expansion of a defining function [^5] A defining function $s\in C^{\infty}(M)$ is defining for $\Sigma$ if $s=0$ and $|ds|_{g}=1$ on $\Sigma$. The question raised in [70] is to find the highest number $\ell\in\mathbb{N}$ and a defining function $s$ such that we have an expansion of the form $|ds|_{g}=1+O(s^{\ell})$ near $\Sigma$. $s$ for $\Sigma$, then there is an obstruction, called the obstruction density. This obstruction density for d1=2d-1=2 (i.e. when $\Sigma$ is a surface) and $M^{d}=\mathbb{R}^{3}$ turns out to be exactly the left-hand side of (I.1). As proved by [72, 68], this obstruction can be understood in general as the Euler–Lagrange equation of a coefficient in some renormalized volume expansion reminiscent to [76, 75]. These alternatives will be discussed in Section II.2.2. One common point of these functionals is that if $\Sigma^{n}\subset\mathbb{R}^{d}$ has even dimension $n\geq 4$, then the Euler–Lagrange equation of any of this generalized Willmore energy has a leading-order term of the form $\Delta_{g_{\Sigma}}^{\frac{n}{2}}\vec{H}_{\Sigma}$. This term is also the leading-order term of the functional
$\vec{\Phi}\in\mathrm{Imm}(\Sigma^{n},\mathbb{R}^{d})\mapsto\int_{\Sigma^{n}}\big|\nabla^{\frac{n-2}{2}}\vec{\mathrm{I\!I}}_{\vec{\Phi}}\big|^{2}_{g_{\vec{\Phi}}}\ d\textup{vol}_{g_{\vec{\Phi}}}.$